There are 5 rows of blocks. In each row there is one more block than the last row. There are 15 blocks right now. How many blocks all together will you have if you have 10 rows?
Answer by stanbon(75887)

There are 5 rows of blocks. In each row there is one more block than the last row. There are 15 blocks right now. How many blocks all together will you have if you have 10 rows?
---
1 row: 1
2 rows: 1+2 = 3
3 rows: 1 + 2 + 3 = 6
4 rows: 1 + 2 + 3 + 4 = 10
5 rows: 1 + 2 + 3 + 4 + 5 = 15
----
Your Problem:
10 rows: 1 + 2 + 3 + 4 + ...+ 8 + 9 + 10
= 5*11 = 55 blocks
=========================
Cheers,
Stan H.